Michelle is eight years older than adam. Three years ago michelle was twice as old as adams was then. How old is Michelle?
-I just need to know how to set it up, very confused on that

Answer by scott8148(6628)

now: m=a+8 ... a=m-8

3 years ago: m-3=2(a-3) ... m-3=2((m-8)-3)